OpenStudy (anonymous):

lim x -->-infinity of sqrt*(x^2 +x +1)+x

OpenStudy (anonymous):

gimmick is to multiply by the conjugate

OpenStudy (anonymous):

i've tried that i can't figure out to do once you get past that

OpenStudy (anonymous):

you end up with x+1/sqrt*(x^2 +x+1) -x

OpenStudy (anonymous):

\[\frac{x+1}{\sqrt{x^2+x+1}-x}\]

OpenStudy (anonymous):

yeah what you wrote

OpenStudy (anonymous):

what i don't get is where to go from there

OpenStudy (anonymous):

Multiply everything by the conjugate again maybe?

OpenStudy (anonymous):

is l'hopital allowed?

OpenStudy (anonymous):

you would end up with \[\sqrt{x^2+x+1}-x^2-x / X^2+1\]

OpenStudy (anonymous):

which i don't think gets me anywhere

OpenStudy (anonymous):

Nvm, that just complicates things.

OpenStudy (anonymous):

Is the answer 0?

OpenStudy (anonymous):

they only thing i've seen is dividing the numerator and denominator by x but I can't seem to do much once i get there the answer is -.5 i just need to show the process

OpenStudy (anonymous):

the answer is \(-\frac{1}{2}\)

OpenStudy (anonymous):

in the numerator that would give me \[ 1+\frac{ 1 }{ x }\] and because it is as x --> negative infinity 1/x equals 0 so the numerator is 1

OpenStudy (anonymous):

but the denominator is what i can't work out

OpenStudy (anonymous):

Yeah I know, so there is another way possibly.

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

you can divide everything by \(x\) just don't bring it inside the radical

OpenStudy (anonymous):

but how does that help then?

OpenStudy (anonymous):

I tried that satellite, I still end up using L'hostipals rules after a while.

OpenStudy (anonymous):

\[\frac{x+1}{\sqrt{x^2+x+1}-x}\] \[=\frac{1+\frac{1}{x}}{-1+\frac{\sqrt{x^2+x+1}}{x}}\]

OpenStudy (anonymous):

Yep I got there.

OpenStudy (anonymous):

I ended up using L'hospital's rule.

OpenStudy (anonymous):

ignore the \(\frac{1}{x}\) up top

OpenStudy (anonymous):

right because that would approach zero

OpenStudy (anonymous):

and \[\lim_{x\to -\infty}\frac{\sqrt{x^2+x+1}}{x}=-1\]

OpenStudy (anonymous):

Wait why is that -1?

OpenStudy (anonymous):

because \(x\) is going to \(-\infty\) so the denominator is negative, while the numerator is positive

OpenStudy (anonymous):

oh and the degress of the terms are equal

OpenStudy (anonymous):

Ohh yeah, right.

OpenStudy (anonymous):

essentially, yes

OpenStudy (anonymous):

Then the rest is simple.

OpenStudy (anonymous):

now you get the \(-\frac{1}{2}\)

OpenStudy (anonymous):

yes thanks, and a really stupid question how do i stop this now that we answerd the question

OpenStudy (anonymous):

Press close.

OpenStudy (anonymous):

ok, thanks